3. If you were to build a pi pe organ with open-open tube pipes that could play notes spanning the entire range of human hearing ), what would be the range of the lengths (L) of pipes? Assume that the fundamental frequency is heard (n the lengths () of p open pipe, A - 2L. See Fig. 17.15 (b) and Equation 17.17. Assume Vound in alr 343 m/s L 20,000 Hz pie L 20 Hz pipe 3B) Would you expect that the longer pipe produces the lower frequency note? Why or why not? Explain.
Answer #1

(A) Speed of sound in air = 343 m/s

Wavelength of 20 Hz wave = 343/20 = 17.15m

Wavelength of 20 kHz wave = 17.15mm

Length of open pipe is approximately half the wavelength (ignoring end corrections)

L (of 20 Hz wave) = 17.15m/2 = 8.575 m

L (of 20 kHz wave) = 8.575 mm

(B) Longer pipe means longer wavelength and frequency is inversely proportional to wavelength. So yes, longer pipe will produce the lower frequency.